# SPADE: Sparsity-Guided Debugging for Deep Neural Networks

### v1

@@ 3.1 Algorithm Overview
10:     end for
11:     for  Each layer in $M$  do
12:         $\tilde{W}_{i}\leftarrow\textnormal{argmin}_{W\textnormal{sparse}}\|WX_{i}-Y_{i}\|_{2}^{2}$
13:         $W_{i}\leftarrow\tilde{W}_{i}$triangleright Replace weights with sparse ones
14:     end for
15:     return $I(M,s)$triangleright Interpretability method on $M,s$
16:end procedure
Figure 2: (Left) The overall SPADE procedure: given an image and a model, SPADE prunes the model using image augmentations. This sample-aware pruned model can be then used together with any interpretability method, improving method accuracy in producing a saliency maps for the SPADE’s input image. (Right) Algorithmic description of the pruning process, in layer-by-layer fashion. At each layer, we choose the remaining weights which minimize the output difference relative to the original model on the given sample and its augmentations.
3.2 Implementation Details
Pruning approach.
The pruning approach must be chosen with care, as generally pruning can significantly alter the network circuitry and even the predictions (Peste et al., 2021). Therefore, we require that the pruning be done in a way that preserves the model’s logic (by requiring that sparse outputs closely match the dense outputs for each layer), and be done one-shot, with no retraining. For this task, one can use one of the existing one-shot sparsity solvers, e.g. (Hubara et al., 2021; Frantar & Alistarh, 2023; 2022; Kuznedelev et al., 2023). We chose the OBC solver (Frantar & Alistarh, 2022), which provides an approximate solution to the $\ell_{2}$ constrained problem in Equation 1.
Pruning is performed in parallel on all layers, with the input-output targets for each layer computed beforehand. Thus, the pruning decisions of each layer are independent of each other. Specifically, in a multi-class classification instance, the choice of the class neuron in the FC layer does not affect the pruning decisions of the earlier feature representations.
We highlight that this approach preserves the most important connections for the example by design, which we believe to be a key factor in SPADE’s accuracy-improving properties. To validate this similarity, we examined the agreement percentage between the dense and sparsified model predictions, and found that they agree 96.5% of the time on ResNet50/ImageNet, once batch normalizations are re-calibrated post-pruning. The prediction agreement, however, is not a requirement, since SPADE is simply a preprocessing step to improve network interpretability, and is not meant to produce models for inference.
Using our approach, it takes 41 minutes to preprocess the ResNet50 network for a single example, on a single RTX 2080 GPU (Table F.15). By comparison, it takes 40 hours to preprocess the network with the FC pruning method of Wong et al. (2021). (However, we note that SPADE must be run once per sample or group of samples, and the FC pruning method is run once for all examples. Irrespective of runtime, experiments in the next section show that our approach is significantly more accurate in practice.) The SPADE runtime may be sped up by only sparsifying the final layers of the network at a small accuracy cost (see Appendix D), and possibly by using more efficient sparsity solvers (Frantar & Alistarh, 2023). We will explore the second direction in future work.
@@ 3.2 Implementation Details
One key question is how to choose the target sparsity ratio to which each layer is pruned, that is, how many weights to remove from each layer. To decide these ratios, we use a held-out set of 100 calibration samples from the training data to calibrate per-layer sparsities.Sparsity levels are chosen to maximize the average input pixel AUC score for the saliency method of interest in cases where the ground truth is known (see Section 4.1). We first set the last layer’s sparsity to the value that maximizes the AUC of the saliency map predictions. Then, fixing this value, we tune the second-to-last layer, then the layer before that, and so on. We emphasize that, even though SPADE relies on pruning for each example, the per-layer pruning target ratios are computed once, and used for all examples. Further, we show in Section D that layer sparsity hyperparameters tuned on ImageNet may be used for other datasets on the same network architecture, and we also present a heuristic-based approach to sparsity ratio tuning that may be used if tuning overhead is a concern in Appendix Section  D.3.
@@
There are two motivations for employing augmentations. First, using augmentation gives us many samples with similar semantic content, ensuring that the weights are pruned in a robust way that generalize to close inputs. Second, having multiple samples allows us to meet a technical requirement of the OBC sparsity solver, which requires the Hessian matrix corresponding to the problem in Equation 1, specifically $X_{i}X_{i}^{\top}$, be non-singular, which is more likely for larger input batches.
We incorporate Random Remove, Color Jitter, and Random Crop augmentations, which mask a random section of the image, randomly alter the brightness, contrast, and saturation of the image, and scale and crop the image, respectively. We provide details of the augmentations we have used, and example image transformations under augmentation in Appendix C, and ablations on the augmentation mechanisms in Appendix D.2.
@@ 4 Experiments
In this section, we experimentally validate the impact of SPADE on the usefulness and the fidelity of network interpretations. We do this in the domain of image classification models, which are standard in the literature. Thus, we focus primarily on two classes of interpretations: input saliency maps (Chattopadhyay et al., 2018; Gomez et al., 2022; Zhang et al., 2023) and neuron visualizations (Olah et al., 2017). Our goals are to demonstrate the following:
@@
    Input saliency maps produced after preprocessing with SPADE accurately identify the image areas responsible for the classification.
@@
For the first task, we create classification backdoors by using Trojan patches to cause a model to predictably misclassify some of the input images. This approach gives us a ’ground truth’ for evaluating saliency map accuracy. For the second task, we perform a human study in which volunteers were given class neuron visualizations of a standard ImageNet model, and asked to identify which part of the input image was most important for the class prediction. Crucially, the ground truth for this study, i.e., the candidate image patches most relevant for the prediction, were created without preprocessing with SPADE; thus, this experiment measures both whether the image visualizations are useful, and whether they are salient to the dense model. Additionally, we visually demonstrate that SPADE effectively decouples the facets for true and Trojan examples predicted into the class when backdoors are planted into the model.
4.1 Impact of SPADE on input saliency map accuracy
@@
We first describe the results of applying SPADE preprocessing before creating saliency maps. Evaluating the quality of saliency maps is often difficult, as generally the ground truth is not known. Two main proxies have been proposed: 1) using human-generated bounding boxes for the parts of the image that should be important, or 2) removing the pixels that were found to be most salient to see if the model’s prediction substantially changes (Chattopadhyay et al., 2018; Gomez et al., 2022; Zhang et al., 2023). Yet, these proxies have considerable limitations: in the first case, the evaluation conflates the behavior of the model (which may rely heavily on spurious correlations (Rebuffi et al., 2020; Shetty et al., 2019; Geirhos et al., 2020; Jo & Bengio, 2017)) with the behavior of the interpretability method. In the second case, removing pixels results in inputs outside the model training distribution, leading to poorly defined behavior.
Therefore, we follow the recent methodology of Casper et al. (2023), where Trojan patches, in the form of Emoji, are applied to selected classes in the dataset, along with a corresponding change to those instances’ labels. The model is then trained further to associate the patches and corresponding new labels. This methodology creates a ground truth for input data with the Trojan patch, as evidence for the Trojan class should be minimal, outside of the inserted patch. Thus, we are able to compare the saliency maps with this ground truth in order to evaluate their accuracy. We use two metrics to assign accuracy scores to saliency maps. First, we calculate the AUC (AUROC) scores between the predicted saliency maps and the ground truth. In this way, the evaluation is not affected by the scale of the saliency map weights but only by their ordering, ensuring that ajdustments don’t need to be made between methods. Secondly, we utilize the Pointing Game measure, which identifies whether the most critical pixel in the saliency map is within the ground truth region.
@@
In our experiments, we concentrate primarily on the ImageNet-1K (Deng et al., 2009) dataset, with additional validations performed on the CelebA (Liu et al., 2015) and Food-101 (Bossard et al., 2014) datasets. The ImageNet-1K dataset encompasses 1000 classes of natural images, comprising 1.2 million training examples.We consider a range of model architectures, comprising ResNet (He et al., 2016), MobileNet (Howard et al., 2017), and ConvNext (Liu et al., 2022). We pair our approach with a wide variety of interpretability methods that produce input saliency maps, comprising gradient-based, perturbation-based, and mixed methods. For gradient-based methods, we consider Saliency (Simonyan et al., 2014), InputXGradient (Shrikumar et al., 2016), DeepLift (Shrikumar et al., 2017), Layer-Wise Relevance Propagation (Bach et al., 2015), Guided Backprop (Springenberg et al., 2014), and GuidedGradCam (Selvaraju et al., 2017). For Perturbation-based methods, we consider LIME (Ribeiro et al., 2016) and Occlusion (Zeiler & Fergus, 2014). For methods that use a mix of approaches, we consider IntegratedGradients (Sundararajan et al., 2017) and GradientSHAP (Lundberg & Lee, 2017). A description of the methods is available in Appendix Section A. We tune sparsity ratios separately for each method used.
Training Details.
We follow Casper et al. (2023) in randomly selecting 400 samples from the ImageNet-1K training set for each Trojan patche. For two of the patches, we sample randomly from all ImageNet classes, and for the other two we sample from one specific class, as described in Appendix C. We then finetune clean pretrained models to plant the backdoors. For experiments on ImageNet, we fine-tune the model using standard SGD-based training for six epochs, with learning rate decay at the third epoch. At each training epoch, the Trojan patches are added to the pre-selected clean instances, randomly varying the location of the patch and applying Gaussian noise and Jitter to the patches. The exact hyper-parameters are provided in Appendix C.
@@
Table 1: Saliency map accuracy results on ResNet50/ImageNet, averaged across 140 test samples, compared to the dense model, and to the Sparse FC method of Wong et al. (2021).
Saliency Method AUC Pointing Game
Dense SPADE Sparse FC Dense SPADE Sparse FC
Saliency 86.92$\pm$7.85 95.32$\pm$7.5 87.19$\pm$7.57 83.92 93.71 81.94
InputXGradient 83.77$\pm$10.21 93.73$\pm$8.59 84.05$\pm$0.95 67.83 88.81 66.67
DeepLift 93.47$\pm$4.17 95.85$\pm$3.92 93.61$\pm$2.42 89.51 90.91 89.58
LRP 90.05$\pm$8.52 99.11$\pm$0.81 93.49$\pm$8.08 72.73 96.5 81.94
GuidedBackprop 95.22$\pm$3.73 96.45$\pm$4.68 95.27$\pm$3.95 87.5 86.81 86.81
GuidedGradCam 97.82$\pm$1.68 98.12$\pm$1.64 97.79$\pm$4.22 90.91 93.71 90.97
LIME 91.93$\pm$8.32 95.84$\pm$3.73 92.57$\pm$9.09 70.63 69.23 70.83
Occlusion 86.09$\pm$11.51 93.73$\pm$9.53 85.79$\pm$24.35 89.51 86.71 88.19
IntegratedGradients 87.86$\pm$8.63 94.77$\pm$8.19 88.33$\pm$1.44 81.12 88.81 83.33
GradientShap 87.74$\pm$8.66 94.85$\pm$7.35 88.23$\pm$1.53 81.12 88.11 81.94
Average 90.09 95.78 90.63 81.41 87.22 82.22
@@
We benchmark our results against the method of Wong et al. (2021), which we will refer to for simplicity as “Sparse FC.” Recall that this method completely retrains the final FC layer via heavy regularization, after which it applies existing interpretability methods. The results on the ImageNet/ResNet50 combination are shown in Table 1. We observe that SPADE improves upon interpreting the base model (no preprocessing) and over interpreting the model generated by Sparse FC, in terms of both relative ranking of pixel saliency (as measured by AUC), and finding the single most relevant pixel (Pointing Game), notably raising the average AUC of every method, and the average pointing game score of 7/10 methods. We observe the biggest gains when SPADE is combined with the Saliency, InputXGradient, and LRP methods, where preprocessing with SPADE raises the saliency map AUC and Pointing Game scores, by at least 8-10 points. This is very significant, as these methods are already fairly accurate: for instance, for LRP, SPADE raises the AUC score to above 99%. On the negative side, while SPADE raises the Pointing Game scores of gradient-based methods, it slightly lowers those scores for the Occlusion and LIME methods, which rely on permutations. The average AUC improvement of our method is 5.69%, whereas the average improvement of SparseFC is 0.54%. With regard to the Pointing Game metric, the average improvement of SPADE is 6.81%, while the Sparse FC method’s average improvement is 0.81%.
Additional validation, and ablation study.
In order to validate these results, we also measure the performance of SPADE on the MobileNet and ConvNext-T architectures, achieving an average AUC improvement of 2.90% for MobileNet and 3.99% for ConvNext. Full results are provided in Appendix B. In addition, we perform an ablation study (see Appendix D) of SPADE’s most salient hyperparameters, with the following conclusions.
First, the layer sparsity targets tuned on the ImageNet dataset transfer well to the CelebA and Food101 datasets. Additionally, it is possible to prune only the final block of the ResNet50 architecture with only a small drop in saliency map accuracy: this reduces the
pruning time and resource usage of the sparsity solver significantly, where pruning speed and resources are a concern. Second, the choice of a single, augmented sample as the pruning dataset far outperforms other options, such as using a random selection from the same class. Finally, SPADE is fairly robust to the specific choice of augmentations applied to the example; however, the best results are obtained when a with a combination of Jitter and Random Crop.
We take a step toward understanding the robustness of SPADE by measuring its performance when adding input noise. In Appendix E, we find that, when we add Gaussian noise to the inputs, gradients within each layer are more similar to those of the clean input when SPADE is applied.

### v2

@@ 3.2 Implementation Details
  end for
  for  Each layer in $M$  do
     $\tilde{W}_{i}\leftarrow\textnormal{argmin}_{W\textnormal{sparse}}\|WX_{i}-Y_{i}\|_{2}^{2}$
     $W_{i}\leftarrow\tilde{W}_{i}$Replace weights with sparse ones}
  end for
  Return $I(M,s)$Interpretability method on $M,s$}
Algorithm 1 SPADE
As an orthogonal contribution, we show that, in our setting, this solver can be sped up significantly by efficiently grouping pruning operations across several inputs on the GPU. With these changes, ResNet50 pruning amortizes to about 3 seconds/example. Going forward, we refer to the versions of SPADE employing the OBS and SparseGPT solvers as SPADE and FastSPADE, respectively. The timings are summarized in Table 1.
Table 1: Per-example timings of different versions of SPADE. Batched FastSPADE is computed on a batch of 25 examples. Timings computed on an NVIDIA GeFORCE GPU with 25GiB RAM.
Pruner type Forward Pass+Hessian Pruning+Saving
SPADE 41s 15m51s
FastSPADE 3s 20s
Batched FastSPADE 1s 2s
Pruning is performed in parallel on all layers, with the input-output targets for each layer computed beforehand. Thus, the pruning decisions of each layer are independent of each other. Specifically, in a multi-class classification instance, the choice of the class neuron in the FC layer does not affect the pruning decisions of other layers. We ablate sequential pruning as an alternative to parallel in Appendix G.1.
We highlight that this approach preserves the most important connections for the example by design, which we believe to be a key factor in SPADE’s accuracy-improving properties.
@@
One key question is how to choose the target layer sparsity ratio, i.e., how many weights to remove from each layer. There are two challenges with tuning the correct sparsity ratios. First, hyperparameter tuning in general may be resource-intensive. Second, we need some measure of ground truth for the saliency method’s correctness. To overcome the first problem, we note that sparsity ratios may be tuned on as few as 100 examples, which is feasible with either version of the method, but especially with FastSPADE. We emphasize that, even though SPADE relies on pruning for each example, the per-layer pruning target ratios are computed once for all examples. Further, we show in Appendix  B that layer sparsity hyperparameters tuned on ImageNet may be used for other datasets on the same network architecture. We also explore a heuristic-based approach to sparsity ratio tuning, as well as experiments showing that it is possible to get improvements using a smaller number of samples, as well as using FastSPADE, in Appendix B.
To overcome the second problem, we propose two approaches. The first is to use Trojan patches in a version of the model that includes backdoors. We validate in Appendix 4.1 that sparsity targets chosen using the Trojan patches method are generally applicable by examining insertion/deletion metrics for pixel attribution on clean input examples, and by using a different set of Trojan patches. Additionally, we show that it is possible to calibrate the layer sparsities using the pixel insertion metric.
For all approaches, sparsity levels are chosen to maximize the desired metric for the saliency method of interest, and tuned in inverse order of layer depth. That is, we first set the last layer’s sparsity to the value that maximizes the metric. Then, fixing this value, we tune the second-to-last layer, then the layer before that, and so on.
@@
There are two motivations for employing augmentations. First, using augmentation gives us many samples with similar semantic content, ensuring that the weights are pruned in a way that generalizes to close inputs. Second, having multiple samples allows us to meet a technical requirement of the sparsity solvers, namely that the Hessian matrix corresponding to the problem in Equation 1, specifically $X_{i}X_{i}^{\top}$, be non-singular, which is more likely for larger input batches. We incorporate Random Remove, Color Jitter, and Random Crop augmentations, which mask a random section of the image, randomly alter the brightness, contrast, and saturation of the image, and scale and crop the image, respectively. We provide details of the augmentations we have used, and example image transformations under augmentation in Appendix F, and ablations on the augmentation mechanisms in Appendix G.3.
@@ 4 Experiments
In this section, we experimentally validate the impact of SPADE on the usefulness and fidelity of network interpretations. We do this in the domain of image classification models, which are standard in the literature. Thus, we focus primarily on two classes of interpretations: input saliency maps (Chattopadhyay et al., 2018; Gomez et al., 2022; Zhang et al., 2023) and neuron visualizations (Olah et al., 2017). Our goals are to demonstrate the following:
@@
    Input saliency maps produced after preprocessing with SPADE accurately identify the image areas responsible for the dense model’s classification.
@@
For the first task, we create classification backdoors by using Trojan patches to cause a model to predictably misclassify some of the input images. This approach gives us a “ground truth” for evaluating saliency map accuracy; we further validate the results by measuring whether the pixels identified by the saliency ranking on clean inputs drive the dense model’s confidence in the prediction. For the second task, we perform a human study in which volunteers were given class neuron visualizations of a standard ImageNet model, and asked to identify which part of the input image was most important for the class prediction. Crucially, the ground truth for this study, i.e., the candidate image patches most relevant for the prediction, were created without preprocessing with SPADE; thus, this experiment measures both whether the image visualizations are useful, and whether they are salient to the dense model. Additionally, we visually demonstrate that SPADE effectively decouples the facets for clean images in Figure 2, and for true and Trojan examples predicted into the class in Appendix J.
4.1 Impact of SPADE on Saliency Map Accuracy
@@
Evaluating the quality of saliency maps is often difficult, as generally the ground truth is not known. Two main proxies have been proposed: 1) using human-generated bounding boxes for the parts of the image that should be important, or 2) inserting or removing the pixels that were found to be most salient to see if the model’s prediction substantially changes (Chattopadhyay et al., 2018; Gomez et al., 2022; Zhang et al., 2023). Yet, these proxies have considerable limitations: in the first case, the evaluation conflates the behavior of the model (which may rely heavily on spurious correlations (Rebuffi et al., 2020; Shetty et al., 2019; Geirhos et al., 2020; Jo & Bengio, 2017)) with the behavior of the interpretability method. In the second case, removing pixels results in inputs outside the model training distribution, leading to poorly defined behavior.
To overcome this issue, a recent paper (Casper et al., 2023) proposed using Trojan patches, in the form of Emoji. These are applied to selected classes in the dataset, along with a corresponding change to those instances’ labels. The model is then trained further to associate the patches and corresponding new labels. This approach creates a ground truth for input data with the Trojan patch, as evidence for the Trojan class should be minimal, outside of the inserted patch. To our knowledge, this is the only approach that enables the comparison of saliency maps with actual ground truth, and so we primarily rely on this method to test the accuracy of SPADE.
We calculate the AUC (AUROC) scores between the predicted saliency maps and the ground truth. In this way, the evaluation is not affected by the scale of the saliency map weights but only by their ordering, ensuring that adjustments don’t need to be made between methods.
We acknowledge, however, that the applicability of this method to other inputs, for instance, images where the evidence for a class may be more dispersed, is not well-understood. We therefore additionally validate SPADE using the Insertion/Deletion metrics introduced by (Petsiuk et al., 2018), which does not rely on Trojan patches. In this evaluation, a saliency method is used to rank all pixels in the image in terms of their relevance to the prediction. These pixels are then either added to a blank image (insertion) or removed from the full image (deletion) in decreasing order of importance, and the AUC(AUROC) score is computed on the confidence (softmax) score of the model for the predicted class, normalized by the softmax score on the full image. We use clean images (without a Trojan patch) for this evaluation, confirming that sparsity targets set using Trojan patches transfer to this use case. Additionally, we use alternate sparsity targets tuned using the Insertion metric, showing that sparsity ratios may be tuned even without having a backdoored model.
@@
We concentrate primarily on the ImageNet-1K (Deng et al., 2009) dataset, with additional validations performed on the CelebA (Liu et al., 2015) and Food-101 (Bossard et al., 2014) datasets. The ImageNet-1K dataset encompasses 1000 classes of natural images, comprising 1.2 million training examples. We consider a range of model architectures, comprising ResNet (He et al., 2016), MobileNet-v2 (Howard et al., 2017), and ConvNext (Liu et al., 2022). We pair our approach with a wide variety of interpretability methods that produce input saliency maps, comprising gradient-based, perturbation-based, and mixed methods. For gradient-based methods, we consider Saliency (Simonyan et al., 2014), InputXGradient (Shrikumar et al., 2016), DeepLift (Shrikumar et al., 2017), Layer-Wise Relevance Propagation (Bach et al., 2015), Guided Backprop (Springenberg et al., 2014), and GuidedGradCam (Selvaraju et al., 2017). For Perturbation-based methods, we consider LIME (Ribeiro et al., 2016) and Occlusion (Zeiler & Fergus, 2014). For methods that use a mix of approaches, we consider IntegratedGradients (Sundararajan et al., 2017a) and GradientSHAP (Lundberg & Lee, 2017). A description of the methods is available in Appendix A. We tune sparsity ratios separately for each method used. We use the Captum library (Kokhlikyan et al., 2020) for saliency method implementations, except for LRP, for which we use (Nam et al., 2019).
@@
Table 2: Saliency map Trojan AUCon ResNet50/ImageNet, averaged across 111 test samples, compared to the dense model, and to the Sparse FC method of Wong et al. (2021).
Saliency Method Dense SPADE FastSPADE Sparse FC
Saliency 87.87 96.21 93.91 88.05
InputXGradient 85.44 95.10 90.61 85.59
DeepLift 94.10 96.55 95.07 94.21
LRP 90.81 99.21 98.03 93.99
GuidedBackprop 95.73 97.08 95.81 95.82
GuidedGradCam 98.03 98.37 97.75 98.00
LIME 90.69 95.47 93.94 91.83
Occlusion 88.29 95.40 90.90 87.84
IntegratedGradients 89.61 96.10 93.55 89.89
GradientShap 89.51 96.03 93.80 89.82
Average 91.01 96.55 94.34 91.50
Table 3: Insertion and Deletion Metric AUCon clean inputs, compared to the dense model, and to the Sparse FC method of Wong et al. (2021). FastSPADE* refers to FastSPADE with layer sparsity targets tuned using the insertion metric.
Saliency Method Insertion $\uparrow$ Deletion $\downarrow$
Dense FastSPADE FastSPADE* Sparse FC Dense FastSPADE FastSPADE* Sparse FC
Saliency 29.90 34.26 33.63 29.78 14.95 12.40 11.33 14.66
InputXGradient 37.36 41.04 34.94 37.61 10.23 8.42 10.17 10.40
DeepLift 42.65 45.26 42.46 43.92 7.78 7.23 5.84 7.80
LRP 46.34 52.92 56.08 52.59 9.71 10.35 11.20 10.68
GuidedBackprop 43.95 43.99 44.77 44.33 9.48 9.70 9.14 9.41
GuidedGradCam 53.90 51.67 54.35 52.83 10.16 9.92 9.89 9.96
LIME 73.19 65.42 68.76 70.20 14.69 16.31 17.82 16.62
Occlusion 32.76 48.95 52.63 32.76 11.58 9.62 10.83 11.65
IntegratedGradients 41.10 43.85 42.09 41.42 8.79 7.03 5.65 8.87
GradientShap 40.62 44.60 42.76 40.98 8.81 7.23 6.26 9.03
Average 44.18 47.20 47.25 44.64 10.62 9.82 9.81 10.91
Backdooring.
For creating Trojan backdoors, we follow Casper et al. (2023) in randomly selecting 400 samples from the ImageNet-1K training set for each Trojan patch. For two of the patches, we sample randomly from all ImageNet classes, and for the other two, we sample from a single class, as described in Appendix F. We then finetune clean pretrained models to plant the backdoors. For experiments on ImageNet, we fine-tune the model using standard SGD-based training for six epochs, with learning rate decay at the third epoch. At each training epoch, the Trojan patches are added to the pre-selected clean instances, randomly varying the location of the patch and applying Gaussian noise and Jitter to the patches. The exact hyper-parameters are provided in Appendix F.
@@
We benchmark our results against the method of (Wong et al., 2021), which we will refer to for simplicity as “Sparse FC.” (Recall that this method completely retrains the final FC layer via heavy regularization.) We use this baseline as it is the closest method to ours in the existing literature and has similar aims; however, note that SPADE is example-specific, while Sparse FC is run globally for all examples. The results on the ImageNet/ResNet50 combination are shown in Table 2. We observe that SPADE improves over using the dense model for interpretation model without preprocessing, and over-interpreting the model generated by Sparse FC, in terms of relative ranking of pixel saliency (as measured by AUC), with SPADE raising the average AUC of every method, and FastSPADE raising the average AUC of 9/10 methods. We observe the biggest gains in Saliency, InputXGradient, and LRP methods, where SPADE raises the saliency map AUC by over 8%, and FastSPADE by over 4%. This is very substantial, as these methods are already fairly accurate: for instance, for LRP, SPADE raises the AUC score to above 99%. However, SPADE produces only small gains for the GuidedBackprop and GuidedGradCam methods, which already have near-perfect accuracy in this study. The average AUC improvement of SPADE is 5.54%, and that of FastSPADE is 3.33%. By comparison, the average improvement of SparseFC is 0.49%.
We present the Insertion Metric results on clean input images for FastSPADE and the Insertion-tuned variant FastSPADE* in Table 3. Preprocessing with FastSPADE improves Insertion and Deletion in 8/10 cases, for an average improvement of 3.02%/9.80%, respectively. FastSPADE* has similar results, improving 8/10 methods on both metrics, and average improvements of 3.07% and 0.81%. The average improvements on the two metrics of Sparse FC, by contrast, are 0.46% and -0.29%.
Additional validation and ablation.
We measure the performance of SPADE on the MobileNet-V2 and ConvNext-T architectures, achieving an average AUC improvement of 2.90% for MobileNet and 3.99% for ConvNext. We also provide initial results of using SPADE with a BERT language model (Devlin et al., 2018), showing gains. Full results are provided in Appendix D. We present an ablation study of SPADE’s most salient hyperparameters in Appendix G.
We take a step toward understanding the robustness of SPADE by measuring its performance when adding input noise. In Appendix I, we find that, when we add Gaussian noise to the inputs, gradients within each layer are more similar to those of the clean input when SPADE is applied.